2x + (9 times sqare root of x) = 5
next question: (square root of x) + square root of (x + 5) = 1

Answer by mukhopadhyay(490)

2x + (9 times sqare root of x) = 5
=> 9[sqrt(x)] = 5-2x
=> [9{sqrt(x)}]^2 = (5-2x)^2
=> 81x = 4x^2-20x+25
=> 4x^2-101x+25 = 0
=> 4x^2-100x-x+25 = 0
=> 4x(x-25)-1(x-25) = 0
=> (x-25)(4x-1) = 0
=> x=25 or x=1/4
...
You can do the other one exactly the same way as this.